In the banana Republic economy, autonomous consumption expenditure is $400 billion, investment is $120 billion, and government expenditure is $90 billion. The marginal propensity to consume is 0.75, and the tax rate is a flat rate of 30% of every $ of every income and autonomous net export is 10 billion. Assumed that the price is fixed in this economy.

a) what is the consumption function?
b) what is the equation of the AE curve?
c) calculate the equilibrium expenditure?
d) calculate the multiplier?
e) if the government run a policy of fiscal discipline where the government expenditure decrease to $30, what is the change in the equilibrium expenditure?
1
Expert's answer
2019-10-14T08:48:42-0400

a) the consumption function is:

C = Ca + c(Y - T) = 400 + 0.75×0.7Y = 400 + 0.525Y.

b) the equation of the AE curve is:

AE = Y = C + I + G + NX.

c) the equilibrium expenditure is:

Y = 400 + 0.525Y + 120 + 90 + 10 = 620 + 0.525Y,

Y = 620/0.475 = 1305.26.

d) the multiplier is:

m = 1/(1 - c(1 - t)) = 1/(1 - 0.75×0.7) = 1/0.475 = 2.11.

e) if the government run a policy of fiscal discipline where the government expenditure decrease to $30, then the change in the equilibrium expenditure is (90 - 30)×2.11 = 126.6.